Solve the equation for the variable:
x^2+2x =63

Answer by faceoff57(108)

x^2+2x =63
x^2 + 2x - 63 = 0
(x + 9)(X - 7) = 0
x + 9 = 0
x = -9
x - 7 = 0
x = 7
x = (-9, 7)